For a first order reaction t0.75 is 138.6sec. it`s specific rate constant is (in s^-1)1. 10^-22. 10^-43. 10^-64. 10^-5

Grade:12

2 Answers

Gopal
6 years ago
For first order reaction  half life and time required for 75% completion are related as  t0.75=2*t0.5     and specific rate constant is related with half life as  k=\tfrac{0.693}{t0.5 }
\therefore t0.5=138.6 /2 =69.25 sec
k=\tfrac{0.693}{69.25 }  ,
k=10 ^{^{-2}}
Ujwal dayma
4 years ago
T0.75=138.6sec
T3/4=138.6sec
2t1/2=138.6sec
T1/2=138.6/2sec
T1/2=69.3
As it is a first order reaction so,
t1/2=0.693/k
Therefore,k=0.693/t1/2
0.693/69.3=
10-2